The electric force between two charges quadruple if the distance between them halves (True,False)
Answer #1

here,

the electric force , F = K * q1 * q2 /r^2

when the distance between them is halved

the electric force F' = K * q1 * q2 /(r/2)^2

F' = 4 * K * q1 * q2 /r^2

F' = 4 F

TRUE , the electric force quadruples